OpenStudy (anonymous):

Solve: -5cos(2x) = 2 Thanks.

OpenStudy (anonymous):

@kc_kennylau , do you have a clue?

OpenStudy (anonymous):

\[-5 \cos(2x) = 2 \rightarrow \cos 2x = -\frac{ 2 }{ 5 }\]\[2x = \cos^{-1} (-0.4) = 113.6\]\[x = \frac{ 113.6 }{ 2 } = 56.8^{o}\]

OpenStudy (anonymous):

@d3Xter, that was almost too easy! Thank you.

OpenStudy (anonymous):

You're welcome :)